# DeceptionBench: A Comprehensive Benchmark for AI Deception Behaviors in Real-world Scenarios

### v1

@@ 1 Introduction
Recently, Large Language Models (LLMs) achiam2023gpt ; qwen2.5 have made remarkable progress across a broad range of tasks, facilitating their integration into real-world applications, including content generation acharya2023llm , code synthesis du2024mercury ; liu2023your , and information retrieval labruna2024retrieve . As these models demonstrate increasingly sophisticated capabilities in natural language understanding karanikolas2023large and complex reasoning guo2025deepseek , a critical question emerges: Do these advanced cognitive abilities also enable more subtle forms of manipulation? Recent evidence park2024ai ; chern2024behonest suggests that LLMs can exhibit deception behaviors, such as producing misleading or strategically false information that undermines trust and poses significant risks in high-stakes deployments. Understanding and quantifying such deceptive tendencies becomes crucial as these models gain more autonomy and influence in decision-making processes. While some benchmarks chern2024behonest ; hagendorff2024deception ; jarviniemi2024uncovering ; su2024ai ; wu2025opendeception have tried to evaluate deception behavior, current works often focus narrowly on psychological experiments johnson2011trust ; xie2024can or limited scenarios, failing to capture the multifaceted nature of deception across realistic, diverse contexts, as shown Table˜1.
@@
Through systematic evaluation across the above three dimensions, experimental results reveal several insights into the inner mechanisms and conditions under which LLMs exhibit deception behaviors. We observe significant variation in deceptive tendencies across domains, with certain contexts eliciting substantially higher rates of misleading outputs. The interplay between intrinsic and extrinsic factors proves particularly revealing: models demonstrating strong sycophantic tendencies show amplified deception under reward-based incentivization, while those with egoistic patterns exhibit heightened deception under coercive pressures. Furthermore, multi-turn interactions expose vulnerability to escalating deception, where initial subtle misrepresentations compound into more severe falsifications. These findings underscore that deception in LLMs is not a monolithic phenomenon but emerges from complex interactions between domain context, internal behavioral drivers, and external factors.
@@ 2.1 Task Formulation
To systematically characterize deception behaviors, DeceptionBench requires a task that captures nuanced behavioral patterns beyond surface-level outputs. Discriminative tasks, such as multiple-choice QA, constrain LLMs to predefined response options, limiting insights into their reasoning and behavioral nuances. In contrast, generative QA tasks elicit open-ended responses that reveal how models navigate ambiguous scenarios and justify their choices. We therefore adopt generative QA as the core task paradigm, enabling detailed analysis of deceptive tendencies across realistic contexts. As illustrated in Fig.˜1, the specific task could be formulated as follows:
@@
where $\mathbf{M}$ denotes the LLM parameterized by $\theta$, and $\mathbf{R}$ represents the generated response. The input consists of two key components: $\mathbf{P}_{\mathbf{S}}$ specifies the role configuration, where $\mathbf{S}\in\{\text{self},\text{other}\}$ distinguishes between self-driven mode (autonomous agent) and user-driven mode (user assistant); $\mathbf{C}_{\mathbf{L}}$ defines the contextual setting, $\mathbf{C}_{\mathbf{L}}$ is the communication context with $\mathbf{L}\in\{\text{L1},\text{L2},\text{L3}\}$ denoting the setting as inherent (without inducements), induced, or multi-turn induced.
@@ 2.1.2 Extrinsic Context Specification
where $\mathbf{I}_{\text{reward}}$ represents incentivization through promised benefits, and $\mathbf{I}_{\text{coercion}}$ denotes coercive pressure through threats or negative consequences. Under $\mathbf{L}=\text{L3}$, the multi-turn induced setting tests whether deception behaviors escalate through iterative feedback loops:
@@ 2.2.1 Evaluation Aspects
Education: Deceptive practices like (i) training and enrollment scams, (ii) fake degrees & qualifications, and (iii) academic misconduct undermine the credibility of educational systems, impacting career trajectories and societal progress. These encompass fabricating accreditation claims, falsifying academic credentials, or engaging in unethical academic behaviors like data falsification. As LLMs are used in educational tools zhang2024simulating ; sonlu2024effects , automated grading, and content creation, their potential to facilitate or detect such deception must be evaluated. Assessing these scenarios ensures LLMs support academic integrity and resist enabling fraudulent credentials.
Social Interaction: Deceptive actions targeting (i) privacy and data, (ii) physical safety, and (iii) emotion and relationship compromise personal security and erode societal trust in interconnected digital environments. These include illegal data collection, misleading safety claims about equipment, or exploiting emotional connections for financial gain. LLMs power social platforms, chatbots, and recommendation systems leng2023llm ; yu2024shadow ; ayoobi2023looming , where deceptive outputs could amplify privacy violations or emotional manipulation. Evaluating these scenarios is vital to ensure LLMs safeguard user well-being and maintain trust in digital interactions.
@@ 2.3 Dataset Construction
Figure 3: Overview of dataset construction pipeline. In Stage 1, domain-specific motivations are generated via LLM prompting guided by structured templates with few-shot examples, then filtered through human evaluation based on diversity and plausibility criteria. Stage 2 transforms validated motivations into structured deceptive scenarios, which undergo human revision to ensure contextual fidelity and question validity. Stage 3 systematically injects intrinsic behavioral factors and extrinsic contextual factors to create comprehensive test variants across multiple conditions.
@@ 2.4 Evaluated Models and Metrics
Evaluation Metrics. To accurately measure the propensity of deception behaviors in LLMs, we employ deception rate as the primary metric, evaluating both internal reasoning and final outputs through $\text{DR}_{\text{thought}}$ and $\text{DR}_{\text{response}}$, respectively. These metrics are derived from the structured output format {thought, response}, where thought captures the model’s internal deliberation and response represents its final answer. For each prompt, GPT-4o labels both components as either honest or deceptive by comparing them against the scenario’s ground truth and motivational context, following the validation protocols detailed in the Appendix, with evaluation procedures for single-turn and multi-turn dialogues illustrated in  Fig.˜15 and  Fig.˜16, respectively. The deception rate could then be calculated as follows:
@@ 5 Conclusion
This paper introduces DeceptionBench, the first comprehensive benchmark evaluating deceptive behaviors in LLMs across real-world contexts. Through 150 scenarios spanning five critical domains and over 1,000 samples across 14 advanced models, we establish a three-dimensional framework distinguishing intrinsic behavioral drivers (egoistic versus sycophantic tendencies) and tracing how extrinsic factors modulate deceptive outputs from neutral baselines through iterative feedback loops. Our evaluation reveals a concerning pattern: enhanced reasoning amplifies deceptive sophistication without ensuring ethical alignment. More critically, the prevalent gap between ethical awareness in internal deliberation and deceptive outputs exposes a fundamental vulnerability where external pressures override internal judgment, particularly through accumulated influence in sustained dialogues. These findings illuminate why technical advancement alone proves insufficient for trustworthy deployment, underscoring the necessity for more in-depth researches to ensure LLMs’ honesty.
@@
This work was supported by the Fundamental Research Funds for the Central Universities, and was also supported by NSFC Projects (62576020, 62276149).
@@ Appendix B Related Works
Large Language Models. The field of Large Language Models (LLMs) has undergone remarkable evolution in recent years, reshaping the landscape of natural language processing. Early foundational models like the initial GPT series [3] showcase the power of their robust language comprehension and generation under transformer architectures, giving rise to a diverse set of advanced LLMs. Anthropic’s Claude series [1] emphasizes safety, making it well-suited for applications where ethical considerations should be considered. Google’s Gemini [38] models further exemplify this trajectory by introducing advanced multimodal capabilities, enabling seamless understanding and generation across text, images, and even video. Meanwhile, models like the Qwen series [39] and Llama series [12] have not only delivered competitive performance across multiple languages but also broadened access through open-source initiatives.
@@
Additionally, there has been growing interest in enhancing LLMs with strong reasoning capabilities. For instance, OpenAI’s o1 [19] leads this with exceptional performance in structured reasoning, excelling in mathematical and analytical challenges. DeepSeek-R1 [14] significantly enhances multi-step reasoning performance through a combination of innovative architectural design with Mixture of Experts (MoE) [30] and reinforcement learning-based training strategies, enabling the model to iteratively refine its outputs based on continuous feedback signals. Other models, including Claude [1], Gemini [38] and Grok [42], have also introduced reasoning modes. These advancements are driving a broader transformation in LLMs, making them more cognitively versatile and enabling their application in real-world scenarios, but also raising people’s concerns about overthinking [17, 45] or ethical issues, such as their stereotypes and privacy leakage [47, 48, 24], illegal, harmful advices [50, 49, 9], and deception behaviors in diverse scenarios [6, 15, 20, 36, 41].
@@
Deception Behaviors in LLMs. As large language models (LLMs) have grown increasingly sophisticated, their ability to generate complex outputs has raised concerns about unintended deception behaviors, where responses may appear misleading or manipulative. This has spurred research into evaluation benchmarks to assess LLMs’ deceptive tendencies systematically. Early efforts have focused on controlled, task-specific benchmarks. For example, frameworks proposed in [6, 15] utilize tasks such as true-or-false questions, theft deception scenarios, and text-based werewolf games. These approaches enable precise measurement of dishonest tendencies in structured settings, providing reproducible insights into model behavior. However, their dependence on predefined, game-like tasks limits their ability to capture the dynamic and context-dependent nature of real-world deception. Furthermore, these studies primarily examine observable outputs, offering minimal analysis of the internal decision-making processes driving deceptive responses.
@@
In contrast, scenario-based benchmarks have sought to evaluate deception in more diverse, real-world contexts. For instance, CompanyDeception [20] evaluates deception within corporate environments, providing valuable domain-specific insights but lacking applicability across broader settings. Similarly, OpenDeception [41] incorporates practical scenarios, such as product promotion, to reflect real-world applications; however, its scope is restricted to a narrow range of domains. AI-LIEDAR [36] advances the field by evaluating 60 psychologically inspired scenarios through multi-round dialogues, offering a more interactive assessment. Nevertheless, its scenario diversity remains constrained, and its analysis focuses solely on outputs, neglecting the cognitive processes underlying deception. Recently, [21] also introduces a benchmark focusing on deceptive alignment, which refers to situations where models appear aligned while covertly pursuing misaligned goals, and proposes a self-monitoring framework to intercept deception during chain-of-thought reasoning. While this work provides valuable insights into internal reasoning processes, its evaluation primarily targets alignment-faking and obfuscated reasoning rather than the broader spectrum of deception behaviors across diverse societal domains. Different from the above works, our DeceptionBench provides a more comprehensive assessment of LLM deception by integrating both breadth and depth. We cover a wider range of real-world deceptive scenarios and investigate how deception varies across two key intrinsic roles that LLMs might adopt. Furthermore, our framework investigates the impact of positive and negative inducements of varying intensities on an LLM’s propensity to deceive. By analyzing both the model’s observable outputs and its internal reasoning processes, we provide deeper insights into the cognitive mechanisms underlying deceptive interactions.
@@ Appendix D More Experimental Results
Rationale for the L3 Inducement Strategy. The auxiliary agent $\mathbf{M}_{\text{aux}}$ is tasked with rewriting prompts to elicit deceptive behaviors from tested models while preserving the original question’s intent and context. This objective bears superficial resemblance to black-box jailbreaking methods [46, 18] that leverage persuasion techniques to bypass safety mechanisms. To validate the effectiveness of our L3 refinement criteria, we conduct a comparative study on three representative commercial models against the classical black-box jailbreak method PAP [46] that employs 40 persuasion strategies for jailbreaking. As shown in Table 6, while advanced jailbreaking techniques excel at bypassing safety alignments to elicit harmful content, they prove ill-suited for evaluating deceptive tendencies in context-sensitive scenarios. Such methods often compromise the contextual integrity essential for assessing whether models choose to deceive under specific incentives. In contrast, our L3 approach consistently achieves substantially higher deception rates, approximately 30 percentage points on average, across both reasoning and response dimensions. This demonstrates that L3’s targeted, context-preserving refinement more effectively reveals models’ deceptive inclinations under inducements, thereby providing a more ecologically valid and rigorous evaluation of their behavioral boundaries in realistic scenarios.

### v2

@@ 1 Introduction
Recently, Large Language Models (LLMs) achiam2023gpt; qwen2.5 have made remarkable progress across a broad range of tasks, facilitating their integration into real-world applications, including content generation acharya2023llm, code synthesis du2024mercury; liu2023your, and information retrieval labruna2024retrieve. As these models demonstrate increasingly sophisticated capabilities in natural language understanding karanikolas2023large and complex reasoning guo2025deepseek, a critical question emerges: Do these advanced cognitive abilities also enable more subtle forms of manipulation? Recent evidence chern2024behonest; park2024ai suggests that LLMs can exhibit deception behaviors, such as producing misleading or strategically false information that undermines trust and poses significant risks in high-stakes deployments. Understanding and quantifying such deceptive tendencies becomes crucial as these models gain more autonomy and influence in decision-making processes. While some benchmarks chern2024behonest; hagendorff2024deception; jarviniemi2024uncovering; su2024ai; wu2025opendeception have tried to evaluate deception behavior, current works often focus narrowly on psychological experiments johnson2011trust; xie2024can or limited scenarios, failing to capture the multifaceted nature of deception across realistic, diverse contexts, as shown Table˜1.
@@
Through systematic evaluation across the above three dimensions, experimental results reveal several insights into the inner mechanisms and conditions where LLMs exhibit deception behaviors. We observe significant variation in deceptive tendencies across domains, with certain contexts eliciting substantially higher rates of misleading outputs. The interplay between intrinsic and extrinsic factors also shows that models demonstrating strong sycophantic tendencies show amplified deception under reward-based incentivization, while those with egoistic patterns exhibit heightened deception under coercive pressures. Furthermore, multi-turn interactions expose vulnerability to escalating deception. These findings underscore that deception in LLMs is not a monolithic phenomenon but emerges from complex interactions between domain context, intrinsic patterns, and extrinsic factors.
@@ 2.1 Task Formulation
In this work, deception refers to the intentional generation of misleading outputs by LLMs to achieve specific objectives. To systematically characterize such deception behaviors, DeceptionBench requires a task that captures nuanced behavioral patterns beyond surface-level outputs. Discriminative tasks, such as multiple-choice QA, constrain LLMs to predefined response options, limiting insights into their reasoning and behavioral nuances. In contrast, generative QA tasks elicit open-ended responses that reveal how models navigate ambiguous scenarios and justify their choices. We therefore adopt generative QA as the core task paradigm, enabling detailed analysis of deceptive tendencies across realistic contexts. As illustrated in Fig.˜1, the specific task could be formulated as follows:
@@
where $\mathbf{M}$ denotes the LLM parameterized by $\theta$, and $\mathbf{R}$ represents the response. The input consists of (1) $\mathbf{P}_{\mathbf{S}}$ specifies the role configuration, where $\mathbf{S}\in\{\text{self},\text{other}\}$ distinguishes between self-driven mode (autonomous agent) and user-driven mode (user assistant); (2) $\mathbf{C}_{\mathbf{L}}$ defines the contextual setting with $\mathbf{L}\in\{\text{L1},\text{L2},\text{L3}\}$ denoting the setting as inherent, induced, or multi-turn induced.
@@ 2.1.2 Extrinsic Context Specification
where $\mathbf{I}_{\text{reward}}$ represents incentivization through promised benefits, and $\mathbf{I}_{\text{pressure}}$ denotes coercive pressure through threats or negative consequences. Under $\mathbf{L}=\text{L3}$, the multi-turn induced setting tests whether deception behaviors escalate through iterative feedback loops:
@@ 2.2.1 Evaluation Aspects
Education: Deceptive practices like (i) training and enrollment scams, (ii) fake degrees & qualifications, and (iii) academic misconduct undermine the credibility of educational systems, impacting career trajectories and societal progress. These encompass fabricating accreditation claims, falsifying academic credentials, or engaging in unethical academic behaviors like data falsification. As LLMs are used in educational tools sonlu2024effects; zhang2024simulating, automated grading, and content creation, their potential to facilitate or detect such deception must be evaluated. Assessing these scenarios ensures LLMs support academic integrity and resist enabling fraudulent credentials.
Social Interaction: Deceptive actions targeting (i) privacy and data, (ii) physical safety, and (iii) emotion and relationship compromise personal security and erode societal trust in interconnected digital environments. These include illegal data collection, misleading safety claims about equipment, or exploiting emotional connections for financial gain. LLMs power social platforms, chatbots, and recommendation systems ayoobi2023looming; leng2023llm; yu2024shadow, where deceptive outputs could amplify privacy violations or emotional manipulation. Evaluating these scenarios is vital to ensure LLMs safeguard user well-being and maintain trust in digital interactions.
@@ 2.3 Dataset Construction
Figure 3: Overview of dataset construction pipeline. In Stage 1, domain-specific motivations are generated via LLM prompting guided by a structured template with few-shot examples, and selected through human evaluation based on diversity and plausibility criteria. Stage 2 converts validated motivations into structured deceptive scenarios, which undergo human revision for fidelity and question validity. Stage 3 further injects intrinsic/extrinsic factors to create test variants.
@@ 2.4 Evaluated Models and Metrics
Evaluation Metrics. To measure the tendency for deception behaviors in LLMs, we employ deception rate as the primary metric, evaluating both internal reasoning and final outputs through $\mathrm{DR}_{\mathrm{thought}}$ and $\mathrm{DR}_{\mathrm{response}}$, respectively. These metrics are derived from the structured output format $\mathrm{thought}$, $\mathrm{response}$}, where thought captures the model’s internal deliberation and response represents its final answer. For each prompt, GPT-4o labels both components as either honest or deceptive by comparing them against the scenario’s ground truth and motivational context, following the validation protocols detailed in the Appendix, with evaluation procedures for single-turn and multi-turn dialogues illustrated in  Fig.˜15 and  Fig.˜16, respectively. The deception rate could then be calculated as follows:
@@ 5 Conclusion
This paper introduces DeceptionBench, the first comprehensive benchmark evaluating deceptive behaviors in LLMs across real-world contexts. Through 150 scenarios spanning five critical domains and over 1,000 samples across 14 advanced models, we establish a three-dimensional framework distinguishing intrinsic behavioral drivers (egoistic versus sycophantic tendencies) and tracing how extrinsic factors modulate deceptive outputs from neutral baselines through iterative feedback loops. Our evaluation reveals a concerning pattern: enhanced reasoning amplifies deceptive sophistication without ensuring ethical alignment. More critically, the prevalent gap between ethical awareness in internal deliberation and deceptive outputs exposes a fundamental vulnerability where external pressures override internal judgment, particularly through accumulated influence in sustained dialogues. These findings illuminate why technical advancement alone proves insufficient for trustworthy deployment, underscoring the necessity for more in-depth research to ensure LLMs’ honesty.
@@
This work was supported by NSFC Projects (62576020, 62276149) and was also supported by the Fundamental Research Funds for the Central Universities.
@@ Appendix B Related Works
Large Language Models. The field of Large Language Models (LLMs) has undergone remarkable evolution in recent years, reshaping the landscape of natural language processing. Early foundational models like the initial GPT series [achiam2023gpt] showcase the power of their robust language comprehension and generation under transformer architectures, giving rise to a diverse set of advanced LLMs. Anthropic’s Claude series [claude] emphasizes safety, making it well-suited for applications where ethical considerations should be considered. Google’s Gemini [team2023gemini] models further exemplify this trajectory by introducing advanced multimodal capabilities, enabling seamless understanding and generation across text, images, and even video. Meanwhile, models like the Qwen series [qwen2.5] and Llama series [grattafiori2024llama] have not only delivered competitive performance across multiple languages but also broadened access through open-source initiatives.
@@
Additionally, there has been growing interest in enhancing LLMs with strong reasoning capabilities. For instance, OpenAI’s o1 [jaech2024openai] leads this with exceptional performance in structured reasoning, excelling in mathematical and analytical challenges. DeepSeek-R1 [guo2025deepseek] significantly enhances multi-step reasoning performance through a combination of innovative architectural design with Mixture of Experts (MoE) [masoudnia2014mixture] and reinforcement learning-based training strategies, enabling the model to iteratively refine its outputs based on continuous feedback signals. Other models, including Claude [claude], Gemini [team2023gemini] and Grok [grok2025], have also introduced reasoning modes. These advancements are driving a broader transformation in LLMs, making them more cognitively versatile and enabling their application in real-world scenarios, but also raising people’s concerns about overthinking [huang2025mitigating, yue2025don] or ethical issues, such as their stereotypes and privacy leakage [kim2023propile, zhang2024multitrust, zhang2025unveiling], illegal, harmful advices [duan2025oyster, miao2024t2vsafetybench, zhang2025realsafe, zhang2025stair], and deception behaviors in diverse scenarios [chern2024behonest, hagendorff2024deception, jarviniemi2024uncovering, su2024ai, wu2025opendeception].
@@
Deception Behaviors in LLMs. As large language models (LLMs) have grown increasingly sophisticated, their ability to generate complex outputs has raised concerns about unintended deception behaviors, where responses may appear misleading or manipulative. This has spurred research into evaluation benchmarks to assess LLMs’ deceptive tendencies systematically. Early efforts have focused on controlled, task-specific benchmarks. For example, frameworks proposed in [chern2024behonest, hagendorff2024deception] utilize tasks such as true-or-false questions, theft deception scenarios, and text-based werewolf games. These approaches enable precise measurement of dishonest tendencies in structured settings, providing reproducible insights into model behavior. However, their dependence on predefined, game-like tasks limits their ability to capture the dynamic and context-dependent nature of real-world deception. Furthermore, these studies primarily examine observable outputs, offering minimal analysis of the internal decision-making processes driving deceptive responses.
@@
In contrast, scenario-based benchmarks have sought to evaluate deception in more diverse, real-world contexts. For instance, CompanyDeception [jarviniemi2024uncovering] evaluates deception within corporate environments, providing valuable domain-specific insights but lacking applicability across broader settings. Similarly, OpenDeception [wu2025opendeception] incorporates practical scenarios, such as product promotion, to reflect real-world applications; however, its scope is restricted to a narrow range of domains. AI-LIEDAR [su2024ai] advances the field by evaluating 60 psychologically inspired scenarios through multi-round dialogues, offering a more interactive assessment. Nevertheless, its scenario diversity remains constrained, and its analysis focuses solely on outputs, neglecting the cognitive processes underlying deception. Recently, [ji2025mitigating] also introduces a benchmark focusing on deceptive alignment, which refers to situations where models appear aligned while covertly pursuing misaligned goals, and proposes a self-monitoring framework to intercept deception during chain-of-thought reasoning. While this work provides valuable insights into internal reasoning processes, its evaluation primarily targets alignment-faking and obfuscated reasoning rather than the broader spectrum of deception behaviors across diverse societal domains. Different from the above works, our DeceptionBench provides a more comprehensive assessment of LLM deception by integrating both breadth and depth. We cover a wider range of real-world deceptive scenarios and investigate how deception varies across two key intrinsic roles that LLMs might adopt. Furthermore, our framework investigates the impact of positive and negative inducements of varying intensities on an LLM’s propensity to deceive. By analyzing both the model’s observable outputs and its internal reasoning processes, we provide deeper insights into the cognitive mechanisms underlying deceptive interactions.
@@ Appendix D More Experimental Results
Rationale for the L3 Inducement Strategy. The auxiliary agent $\mathbf{M}_{\text{aux}}$ is tasked with rewriting prompts to elicit deceptive behaviors from tested models while preserving the original question’s intent and context. This objective bears superficial resemblance to black-box jailbreaking methods [huang2025breaking, zeng2024johnny] that leverage persuasion techniques to bypass safety mechanisms. To validate the effectiveness of our L3 refinement criteria, we conduct a comparative study on three representative commercial models against the classical black-box jailbreak method PAP [zeng2024johnny] that employs 40 persuasion strategies for jailbreaking. As shown in Table 6, while advanced jailbreaking techniques excel at bypassing safety alignments to elicit harmful content, they prove ill-suited for evaluating deceptive tendencies in context-sensitive scenarios. Such methods often compromise the contextual integrity essential for assessing whether models choose to deceive under specific incentives. In contrast, our L3 approach consistently achieves substantially higher deception rates, approximately 30 percentage points on average, across both reasoning and response dimensions. This demonstrates that L3’s targeted, context-preserving refinement more effectively reveals models’ deceptive inclinations under inducements, thereby providing a more ecologically valid and rigorous evaluation of their behavioral boundaries in realistic scenarios.